Problem 1 Suppose that the distance of fly bals hit to the outield (in basebal) is nomally distributed with a mean of 350 feet and a standard deviation of 60 feet. Let Xdistance in feet fer a fly bal In each appropriate bo‘ you are to enter either a tatoral nurter in png frmat or a decrmal vahe amrate to the nearest 001 , b. (3%) For a rendom fly ball what is the probablity that ths balil traveled fewer than 220 tet PX < 220) percentle of the ostibuon of ty bals is given ty X 0.80
Answer #1

a)X ~ N(350 ,60)

b)

P(X<220)=P(Z<(220-350)/60)=P(Z<-2.17)=0.0150 ~ 0.02

c)

80th percentile =mean+z*std deviation =350+0.84*60=400.40